# One missing piece in Vision and Language: A Survey on Comics Understanding

### v1

@@ 4.1 Annotations overview
The available current comics datasets were introduced for different tasks. They vary in the year of release, the language and the origin. This is shown in Table II. From the table, it is noticeable that the available datasets vary significantly in annotation types, making them cover different tasks.
@@
eBDtheque [57] offers comprehensive annotations in high quality for Character, Balloon, and Panel detection, along with Text-Character association. This makes it usable for object detection and linking tasks but it comes with a drawback: it is composed of only 100 pages. In contrast, the COMICS dataset [8] focuses on a set of tasks called closure, based on automatic panel and text detection over a large number of pages. However, it was introduced with poor annotations. Only recently some works [70] have improved OCR text detection on the COMICS dataset. In the same vein, [72] shows that improving OCR accuracy (using Textract) could lead to marginal accuracy gain in closure tasks for the COMICS dataset. Both [70, 72] released the novel OCR annotations to the research community.
It is to note also that some of the dataset with high-quality annotations and large number of pages are not accessible. For instance, BCBId [69], though promising high-quality annotations, falls short in delivery. Similarly, VLRC [47], while rich in annotations, lacks digital applicability due to its focus on physical books.
Manga109 [61] is the first to provide high-level annotations, covering more than 20k pages. The dataset has been improved over several iterations, further expanding the scope of the dataset by adding annotations for onomatopoeias [4] and dialog [73]. The actual state of Manga109 is the result of many years of annotation and improvement by subsequent developments. More recently, following the Manga109 wave, PopManga [41] has been proposed as an English-language alternative, including annotations for recognition and linking, but on a smaller scale (2k images).
@@
In summary, while datasets for Comics Understanding exist, there is a need for more approaches to standardize and unify their annotations. The Comics Dataset Framework [2] proposed a unification of detection annotations as well as a standardization of metrics and settings used to train and measure models. The work was subsequently extended in CoMix [42], where mainly English-language comics (and manga) were selected and the annotations extended to the task of single-page dialogue transcription. In this work, performances of several state-of-the-art models, including Magi [41], CLIP [14], DINOv2 [39], and GPT4 [74], were measured, for the first time.
@@
However, as highlighted in the first column of Table II, the complexity of annotations and tasks is limited. Object detection (panel, characters, texts), text-to-character linking, and re-identifications are integral tasks to understanding the narrative and dialogue in comics. These tasks, along with more recent developments, will be examined in the context of our taxonomy in the following Section.
4.2 Datasets overview
The domain of Comics Understanding encompasses a range of datasets focused on processing and examining comic media. A review of various datasets, as detailed in Table II, reveals that several are no longer accessible (marked with ) or have restricted access (). Those available () often require approval from respective research groups in order to obtain them. This is a common practice in the field of “sharing copyrighted materials” as in fact they can be used only for research purposes in the majority of cases. If the copyright does not allow for redistribution, it is a common practice to release annotations and a script to download and structure the dataset accordingly (as for [41]). As we can see from Table II, many datasets share common origins (column “From”). Notably, these datasets vary significantly in terms of style, quality, and size. The common languages of these datasets are English, French, and Japanese and none of the available data seems to be collected from recent books (after 2010), probably due to copyright limits. We found the above limitation to be one of the main issues in the comic domain: (almost) every work proposes its own (sometimes private) dataset on which results are drawn with little comparison, making these works hard to replicate. Datasets are limited in size and the model’s code and weights are rarely shared. Luckily, this has started changing recently [2, 41, 42].
TABLE III: Number of single instance annotation (declared) in papers. The “Obj. cls” stands for Object-based classification, and refers to the number of classes available if one wants to classify the detected instances in fine-grained classes.
Mode Name Panel Char. Face Textbox Onomat. Obj. cls
Color Fahad18 [63] - 18 - - - 18
Color Ho42 [71] 200 100 - - - -
Color eBDtheque [57] 850 1550 - 1092 - 5
Color SSGCI [66] - 50 - - - 1
Color COMICS [8, 70] 1229664 - - 2498657 - 2
BW Comics3w [75] 159529 - - - - 1
BW JC2463 [34] - - 15801 - - 2
Color AEC912 [34] - - 8184 - - 2
Color DCM772 [58] 4470 10757 5438 - - 4
BW Manga109 [61, 35, 4] 103900 157152 118715 147918 61465 4
Color EmoRecCom [62] 10199 - - - - 8
BW PopManga [41] - 18783 - 20843 - 2
mix CoMix [42] 22176 48903 32625 37923 - 4
In Table III we report, for every dataset, the declared numbers of annotations available. On the right-most column, we report the number of classes annotated, which correspond to different class objects for the detection, or the number of classes for classification annotations. For instance, in Fahad18 there are 18 characters, thus the number of classes is 18. However, in eBDtheque, there are 5 classes, but only panel, character, and face annotations. This is because the characters are separated into three sub-categories: human-like, animal-like, and object-like. Here, we describe the datasets chronologically.
eBDtheque: The eBDtheque [^1] http://ebdtheque.univ-lr.fr/registration [57] is a collection of 100 pages over 20 books, selected from J-comi222previously http://www.j-comi.jp, now no longer available., mainly in the French language (with also English and Japanese comics). [76] led a huge effort in labeling 850 panels, 1092 balloons, 1550 characters, and 4691 text lines, provided by domain experts. For balloons, they provide instance segmentation and classification ( speech, thought, narration, illustrative).
Manga109: The Manga109333http://www.manga109.org/index_en.php [61] dataset contains 109 manga volumes from 93 different authors. Authors successfully obtained the copyrighted material from manga drawers with a “restrictive term-of-use” which implies redistribution only with the scope of research and the possibility to use them in publication with proper referencing the author. These mangas were published between the 1970s and 2010s and are categorized into 12 different genres such as fantasy, humor, sports, etc. This dataset has been further extended by COO [4] which added onomatopoeias polygons annotations and links among truncated onomatopoeias. In fact, on average there are 5.8 onomatopoeias per page in Manga109, while COMICS has much fewer (0.09). Another set of annotations comes from Manga-Dialog [73] which added a text-character link for every speaking text box.
COMICS: The COMICS [^4] https://obj.umiacs.umd.edu/comics/index.html [8] contains 1,229,664 panels paired with automatic textbox transcriptions using an old-OCR system. The 3,948 books are gathered from the Digital Comics Museum (American golden-age comics books) published between 1938 and 1954. The dataset includes ground truth (manually annotated) labeled data such as the rectangular bounding boxes of panels on 500 pages and 1,500 textboxes. Subsequent works have expanded original annotations with better OCR transcriptions [70, 72]. In particular, [70] developed a pipeline for OCR processing and labeling of comic books from COMICS, while [72] extracted all OCR transcripts from COMIC using new-generation OCR textract [^5] Used through API at https://docs.aws.amazon.com/textract/. Lastly, [77] includes strips of dialog prediction based on “persona” information.
DCM772: The DCM772666https://git.univ-lr.fr/crigau02/dcm_dataset [58] contains 772 (pages) images from 27 golden-age comic books from the same COMICS source. Annotations, however, are more precise as they contain panels, the associated characters bounding boxes, and also face boxes.
@@ 4.2 Datasets overview
Figure 4: Example of different comic datasets, from Black-and-White to Color, from comics-style to Manga-style.
4.3 Datasets for Evaluation
@@
Across the datasets presented in Table II, two recent datasets are provided as evaluation sets. In particular, PopManga proposes two splits (validation and test) while CoMix enhances annotations for various existing datasets among which also PopManga.
@@
PopManga: The PopManga [^7] https://github.com/ragavsachdeva/Magi [41] dataset contains English manga titles from the most popular mangas. The dataset contains two test splits: test-seen with 1,136 pages and test-unseen with 789 pages, obtained from 15 and 10 books respectively. The test sets have been annotated with Text (both textboxes and onomatopoeias annotated as text), and Characters.
@@
CoMix: The CoMix [^8] https://github.com/emanuelevivoli/comix-dataset [42] dataset contains English and French comics and manga titles from the most popular datasets: eBDtheque, COMICS, DCM, and PopManga. The dataset contains two splits: validation (available with annotations) and the held-out test split (available only through the evaluation server [^9] Accessible at Robust Reading Competition website). It contains 3.8k images across 100 books, and annotations for multi-task, from detection and linking to dialog generation.
5 Taxonomy: LoCU
@@
Relevant Taxonomies. Initial inspiration stems from the breadth and depth of multimodal questions covered in recent benchmarks like MMMU [78], highlighting the necessity for a taxonomy that accounts for modal complexity and task-specific intricacies in comics. This need is further echoed in the domain-specific taxonomy proposed for video understanding in [79], particularly their dimensional analysis approach, which informs our consideration of spatial and temporal domains in comics. Shifting focus to VL multimodal evaluation, [80] reframes VL tasks as Visual Question Answering (VQA), emphasizing a range of abilities from grounding to numerical calculation. While this framework offers valuable categorizations, it lacks a comprehensive structure for VL tasks, particularly those pertinent to comics. Similarly, the multimodal challenges outlined in [81] offer insights into the representation, translation, alignment, fusion, and co-learning challenges in multimodal learning. These challenges guide our approach to defining taxonomic classes and subclasses within the VL domain, though with a more task-centric focus. Drawing further on [82], we observed the classification of VL tasks into grounding, retrieval, understanding, and generation, with specific adjustments and additions to accommodate comic-specific tasks. Our taxonomy expands on this structure, adding categories such as tagging, augmentation, analysis, segmentation, modification, and synthesis, addressing the wider range of tasks observed in comic studies.
In Table I, we present 31 Vision-Language multimodal tasks, categorized according to the modalities involved, forming the backbone of the LoCU framework. This taxonomy, spanning 10 distinct task groups across five layers, aims to capture the multifaceted nature of Comics Understanding. Layer 0, representing fundamental image processing and viewing capabilities, sets the stage for increasingly complex interactions in the subsequent layers. More details are provided in Appendix B. Taxonomy. This structured approach allows for a comprehensive and nuanced understanding of the range and complexity of tasks within the comics domain, laying a foundation for future exploration and innovation in comic analysis and understanding.
6 Layer 1: Tagging and Augmentation
@@
6.1 Tagging
Tagging tasks predominantly involve classification outputs, albeit with varying input types.
6.1.1 Image Classification
Definition: The task of Image Classification requires classifying an image into one of several predefined categories. In the comics domain, the task can be applied at panel or page level. In both cases, the classes vary from artist names, comic/manga styles or the image type (whether it’s an ads or front page or story page).
Early efforts in this domain, such as those by Chu et al. [31], focused on classifying manga panels by artistic style using Support Vector Machines (SVM) and manga-specific feature vectors. Their feature vectors are built on edge-detected lines creating a 20-dimensional vector from elements such as angle, orientation, the density of segments, etc. Hiroe et al. [83] employed a novel method using SVMs to classify comic books based on the frequency of exclamation marks, demonstrating marks-per-page histogram as an innovative approach to content approximation.
Daiku et al. [84] expanded the scope to page-level classification within manga stories using Convolutional Neural Networks (CNN), distinct from Page Stream Segmentation (discussed in Subsection 6.1.4). Jiang et al. [85] introduced the Consensus Style Centralizing Auto-Encoder (CSCAE) for style classification, particularly distinguishing between Shonen (for boy) and Shojo (for girl) manga styles using robust style feature representations. They used low-rank and group sparsity constraints for consensus for the classification of manga character faces. Lastly, in [86] Terauchi et al. presented an approach for classifying manga based on the author’s unique style, employing a Variational Autoencoder (VAE) for this purpose. They also proposed a “four-scene comics story dataset” for this classification task into Mow, Seinen, and Shonen touches (styles).
In a more recent work [87], the authors tackled the challenge of (multi-)genre classification at the story level in Manga. They introduced a Panel-Page-aware Comic genre classification model, which takes page sequences as input and produces class-wise probabilities, relying on the attention mechanism to jointly attend page features and panel boxes, then pooled into a transformer encoder. The transformer classification token is merged with a Graph Convolution Network (GCN) over the labels graph, which predicts the probability distribution of the labels for the comic stories.
6.1.2 Emotion Classification
@@
A seminal study by Tanaka et al. [88] explored the interplay between speech balloon shapes and emotions in comics. They analyzed the relationship between balloon styles and the emotions expressed within them, using the Manga109 dataset as a foundation for their study.
The EmoReCon challenge at ICDAR 2021 [62] marked a significant advancement in this field. Participants were tasked with classifying 8 emotions from comic panels, using a dataset specifically designed for this purpose [62]. Notably, one of the winning teams employed a three-level fusion approach (early, mid, and late), integrating EfficientNetB3 [89] and RoBERTa [90] for visual and textual embeddings, respectively. This mid-fusion strategy outperformed other architectural configurations, signifying the potential of multimodal fusion. The second winning team proposed an early fusion approach with ResNet50 image features and OCR-extracted text as input for a transformer encoder BERT-style. In this case, the prediction is drawn from a weighted average over the layers’ cls tokens, with comparable results with the previous method.
Not only have the emotions of comic characters been analyzed, but in [91] authors studied readers’ emotions through physiological signals. Through the relationship between a change in the reader’s physiological signal and the page of the reading, authors classify the manga across three classes: comedy, romance, and horror.
A recent work [92] introduced two noteworthy datasets derived from Manga109: MangaAD+ and MangaEmo+. While MangaAD+ is a curated collection of 600 pages featuring unannotated elements such as shop names and ads - in Manga109 only the balloon text and onomatopoeias are annotated -, MangaEmo+ extends this dataset by annotating each page with a range of 8 emotions for multi-label classification. They employed a modified version of Faster R-CNN, trained on augmented text data, for atypical text detection. The subsequent emotion classification task was approached using EfficientNet, which extracted visual features combined with the above custom text detection model. However, neither of the two datasets has been made publicly available.
6.1.3 Action Detection

### v2

@@ 4.1. Annotations overview
The available current comics datasets were introduced for different tasks. From Table 2 it is noticeable that the available datasets vary significantly in size, languages, origin, and -more importantly- annotations, making them cover different tasks.
@@
eBDtheque (Guérin et al., 2013) offers comprehensive annotations in high quality for Character, Balloon, and Panel detection, along with Text-Character association. This makes it usable for object detection and linking tasks but it comes with a drawback: it is composed of only 100 pages. In contrast, the COMICS dataset (Iyyer et al., 2016) focuses on a set of tasks called closure, based on automatic panel and text detection over a large number of pages. However, it was introduced with poor annotations. Only recently some works (Soykan et al., 2022) have improved OCR text detection on the COMICS dataset. In the same vein, (Vivoli et al., 2024d) shows that improving OCR accuracy (using Textract) could lead to marginal accuracy gain in closure tasks for the COMICS dataset. Both (Soykan et al., 2022; Vivoli et al., 2024d) released the novel OCR annotations to the research community. It is to note also that some of the dataset with high-quality annotations and large number of pages are not accessible. For instance, BCBId (Dutta et al., 2022), though promising high-quality annotations, falls short in delivery. Similarly, VLRC (Cohn, 2023), while rich in annotations, lacks digital applicability due to its focus on physical books. Manga109 (Fujimoto et al., 2016) is the first to provide high-level annotations, covering more than 20k pages. The dataset has been improved over several iterations, further expanding the scope of the dataset by adding annotations for onomatopoeias (Baek et al., 2022) and dialog (Li et al., 2023a). The current state of Manga109 is the result of many years of annotation and improvement by subsequent developments. More recently, following the Manga109 wave, PopManga (Sachdeva and Zisserman, 2024) has been proposed as an English-language alternative, including annotations for recognition and linking, but on a smaller scale (2k images).
@@
In summary, while datasets for Comics Understanding exist, there is a need for more approaches to standardize and unify their annotations. The Comics Dataset Framework (Vivoli et al., 2024c) proposed a unification of detection annotations as well as a standardization of metrics and settings used to train and measure models. The work was subsequently extended in CoMix (Vivoli et al., 2024a), where mainly English-language comics (and manga) were selected and the annotations extended to the task of single-page dialogue transcription. In this work, performances of several state-of-the-art models, including Magi (Sachdeva and Zisserman, 2024), CLIP (Radford et al., 2021), DINOv2 (Oquab et al., 2023), and GPT4 (OpenAI, 2023), were measured, across various datasets, for the first time.
@@
However, as highlighted in the first column of Table 2, the complexity of annotations and tasks is limited. While object detection and linking are integral tasks to understanding the narrative and dialogue in comics, we must advance toward more complex tasks found in Vision-Language research. These will be discussed within our taxonomy in the next section.
4.2. Datasets overview
The domain of Comics Understanding encompasses a range of datasets focused on processing and examining comic media. A review of various datasets, as detailed in Table 2, reveals that several are no longer accessible (marked with ) or have restricted access (). Those available ( ✓) often require
@@ 4.2. Datasets overview
Figure 4. Example of different comic datasets, from Black-and-White to Color, from comics-style to Manga-style.
approval from respective research groups to obtain them. This is a common practice in the field of “sharing copyrighted materials” as in fact they can be used only for research purposes in the majority of cases. If the copyright does not allow for redistribution, it is a common practice to release annotations and a script to download and structure the dataset accordingly (as for (Sachdeva and Zisserman, 2024)). As we can see from Table 2, many datasets share common origins (column “From”). Notably, these datasets vary significantly in terms of style, quality, and size. The common languages of these datasets are English, French, and Japanese and none of the available data seems to be collected from recent books (after 2010), probably due to copyright limits. We found the above limitation to be one of the main issues in the comic domain: (almost) every work proposes its own (sometimes private) dataset on which results are drawn with little comparison, making these works hard to replicate. Datasets are limited in size and the model’s code and weights are rarely shared. Luckily, this has started changing recently (Vivoli et al., 2024c; Sachdeva and Zisserman, 2024; Vivoli et al., 2024a).
[…]
Mode / Name Panel Char. Face Text Cls
[…]
In Table 3 we report, for every dataset, the declared numbers of annotations available. On the right-most column, we report the number of classes annotated, which correspond to different class objects for the detection, or the number of classes for classification annotations. For instance, in Fahad18 there are 18 characters, thus the number of classes is 18. However, in eBDtheque, there are 5 classes, but only panel, character, and face annotations. This is because the characters are separated into three sub-categories: human-like, animal-like, and object-like. Here, we describe the datasets chronologically and provide an overview of the image style in Figure 4.
eBDtheque: The eBDtheque [^1] http://ebdtheque.univ-lr.fr/registration (Guérin et al., 2013) includes 100 pages from 20 books, mainly in French, with some English and Japanese comics. (Rigaud, 2014) labeled 850 panels, 1,092 balloons, 1,550 characters, and 4,691 text lines, offering expert annotations. Balloon labels include instance segmentation and classification (speech, thought, narration, illustrative).
Manga109: The Manga109222http://www.manga109.org/index_en.php (Fujimoto et al., 2016) dataset includes 109 manga volumes from 93 authors, with copyrighted material obtained under a “restrictive term-of-use”, allowing use for research and publications with proper citation. The mangas, published between the 1970s and 2010s, span 12 genres. The dataset was extended by COO (Baek et al., 2022), which added onomatopoeia annotations (5.8 per page on average), and Manga-Dialog (Li et al., 2023a), which linked text boxes to characters.
COMICS: The COMICS [^3] https://obj.umiacs.umd.edu/comics/index.html (Iyyer et al., 2016) dataset contains 1,229,664 panels with automatic textbox transcriptions from golden-age American comics (1938-1954). It includes 3,948 books from the Digital Comics Museum and manually annotated bounding boxes for 500 pages and 1,500 textboxes. Later works improved OCR transcriptions (Soykan et al., 2022; Vivoli et al., 2024d), with pipelines developed for better OCR processing and transcript extraction. (Agrawal et al., 2023) added dialog prediction based on ”persona” information.
DCM772: The DCM772444https://git.univ-lr.fr/crigau02/dcm_dataset (Nguyen et al., 2018) contains 772 (pages) images from 27 golden-age comic books from the same COMICS source. Annotations, however, are more precise as they contain panels, the associated characters bounding boxes, and also face boxes.
4.3. Datasets for Evaluation
@@
Across the datasets presented in Table 2, two recent datasets are provided as evaluation sets. In particular, PopManga proposes two splits (validation and test) while CoMix enhances annotations for various existing datasets among which also PopManga.
@@
PopManga: The PopManga [^5] https://github.com/ragavsachdeva/Magi (Sachdeva and Zisserman, 2024) dataset contains English manga titles from the most popular mangas. The dataset contains two test splits: test-seen with 1,136 pages and test-unseen with 789 pages, obtained from 15 and 10 books respectively. The test sets have been annotated with Text (both textboxes and onomatopoeias annotated as text), and Characters.
@@
CoMix: The CoMix [^6] https://emanuelevivoli.github.io/CoMix-dataset (Vivoli et al., 2024a) dataset contains English and French comics and manga titles from the most popular datasets: eBDtheque, COMICS, DCM, and PopManga. The dataset contains two splits: validation (available with annotations) and the held-out test split (available only through the evaluation server [^7] Accessible at Robust Reading Competition website). It contains 3.8k images across 100 books, and annotations for multi-task, from detection and linking to dialog generation.
5. Taxonomy: LoCU
@@
Relevant Taxonomies. Inspiration for this taxonomy comes from multimodal benchmarks like MMMU (Yue et al., 2023), which highlight the need for a structure addressing modal complexity and task-specific intricacies in comics. The dimensional analysis in video understanding from (Fan et al., 2024) also informs our spatial and temporal considerations for comics. In VL evaluation, (Fu et al., 2023) reframes tasks as Visual Question Answering (VQA), but lacks a comprehensive structure for comics-related tasks. Challenges in multimodal learning from (Baltrušaitis et al., 2017) guide our classification of VL tasks, with a task-centric focus. (Li et al., 2022) provides a foundational classification of VL tasks into grounding, retrieval, understanding, and generation. Their work is crucial in shaping our approach, and we build upon their framework by extending these categories to account for comic-specific tasks. In particular, we add categories like tagging, augmentation, analysis, segmentation, modification, and synthesis, which better capture the diversity of tasks observed in comic studies and expand the structure to fully represent the unique challenges in comics understanding.
In Table 1, we present 31 Vision-Language multimodal tasks, categorized according to the modalities involved, forming the backbone of the LoCU framework. This taxonomy, spanning 10 distinct task groups across five layers, aims to capture the multifaceted nature of Comics Understanding. Layer 0, representing fundamental image processing and viewing capabilities, sets the stage for increasingly complex interactions in the subsequent layers. More details are provided in Supplementary Materials. This structured approach allows for a comprehensive and nuanced understanding of the range and complexity of tasks within the comics domain, laying a foundation for future exploration and innovation in comic analysis and understanding.
6. Layer 1: Tagging and Augmentation
@@
6.1. Tagging
Figure 5. Illustration of Tagging tasks.
Tagging tasks predominantly involve classification outputs, albeit with varying input types (Fig. 5).
6.1.1. Image Classification
Definition: The task of Image Classification requires classifying an image into one of several predefined categories. In the domain of comics, the task can be applied at the panel or page level. In both cases, the classes vary from artist names, comic/manga styles, or image type (whether it’s an ad, a front page, or a story page).
Early efforts in this domain, such as those by Chu et al. (Chu and Chao, 2014), focused on classifying manga panels by artistic style using Support Vector Machines (SVM) and manga-specific feature vectors. Their feature vectors are built on edge-detected lines creating a 20-dimensional vector from elements such as angle, orientation, the density of segments, etc. Hiroe et al. (Hiroe and Hotta, 2017) employed a novel method using SVMs to classify comic books based on the frequency of exclamation marks, demonstrating marks-per-page histogram as an innovative approach to content approximation. Daiku et al. (Daiku et al., 2017) expanded the scope to page-level classification within manga stories using Convolutional Neural Networks (CNN), distinct from Page Stream Segmentation (discussed in Subsection 6.1.4). Jiang et al. (Jiang et al., 2018) introduced the Consensus Style Centralizing Auto-Encoder (CSCAE) for style classification, particularly distinguishing between Shonen (for boy) and Shojo (for girl) manga styles using robust style feature representations. They used low-rank and group sparsity constraints for consensus for the classification of manga character faces. Lastly, in (Terauchi et al., 2019) Terauchi et al. presented an approach for classifying manga based on the author’s unique style, employing a Variational Autoencoder (VAE) for this purpose. They also proposed a “four-scene comics story dataset” for this classification task into Mow, Seinen, and Shonen touches (styles). In a more recent work (Xu et al., 2023), the authors tackled the challenge of (multi-)genre classification at the story level in Manga, introducing a Panel-Page-aware Comic genre classification model that takes page sequences as input and produces class-wise probabilities. Their proposal relyied on the attention mechanism to jointly attend page features and panel boxes, then pooled into a transformer encoder. The transformer classification token is merged with a Graph Convolution Network (GCN) over the labels graph, which predicts the probability distribution of the labels for the comic stories.
6.1.2. Emotion Classification
@@
Tanaka et al. (Tanaka et al., 2015) conducted a seminal study on the relationship between speech balloon shapes and emotions in comics, using the Manga109 dataset to analyze how different balloon styles convey emotions. The EmoReCon challenge at ICDAR 2021 (Nguyen et al., 2021b) pushed the field forward by tasking participants with classifying 8 emotions from comic panels using a specially designed dataset. One winning team employed a three-level fusion strategy (early, mid, and late), combining EfficientNetB3 (Tan and Le, 2019) for visual embeddings and RoBERTa (Liu et al., 2019) for textual embeddings. Their mid-fusion approach outperformed other methods, showcasing the power of multimodal fusion. Another top team used an early fusion method, integrating ResNet50 image features with OCR-extracted text processed by a transformer encoder. They predicted emotions using a weighted average of cls tokens, achieving results comparable to the previous team. Beyond character emotions, Sanches et al. (Sanches et al., 2016) analyzed readers’ emotions by tracking physiological signals. By monitoring changes in the reader’s physiological responses, they classified manga into three genres: comedy, romance, and horror. Yang et al. (Yang and Chu, 2023) introduced two new datasets, MangaAD+ and MangaEmo+, based on Manga109. MangaAD+ adds unannotated elements like shop names and ads, while MangaEmo+ includes annotations for 8 emotions on each page for multi-label classification. Using a modified Faster R-CNN for text detection and EfficientNet for visual features, they tackled the emotion classification task, though neither dataset has been made publicly available.
6.1.3. Action Detection